Unstretched length of the bungee cord is L=15 m. Bungee-jumper has a mass of 60 kg. Stiffness of the cord is k=300 N/m. Determine the lowest point that the jumper would reach, assuming that his/her initial velocity is zero. The weight of the cord is neglected.Center of gravity of the jumper is 1.0 m. above the ground
1
Expert's answer
2019-01-15T08:59:53-0500

Since the system is conservative (only potential energies of the cord and the jumper are involved) we can write law of conservation of energy using

"h"

- height of the center of gravity above the ground,

"x"

- the elongation of the spring, and assuming that

"g=10"

:

"mg(L+h+x)=\frac{kx^2}{2},""x^2 \cdot \frac{k}{2}-x\cdot mg-mg(h+L)=0,""x^2 \cdot \frac{300}{2}-x\cdot 60\cdot 10-60\cdot 10(1+15)=0,"

solve this equation and choose the correct root:

"x=10.24 \text { m}."

Thus the lowest point is

"P=L+x=15+10.24=25.24 \text{ m}"

below the surface of the bridge.